# Capacity Approximations for Insertion Channels with Small Insertion Probabilities

### v3

@@ IV Capacity Approximation for the Simple Insertion Channel Model
Let $C(\alpha)$ be the capacity of the simple insertion channel with insertion probability $\alpha$. Then, for small $\alpha$ and any $\epsilon>0$, we have
$C(\alpha)=1+\alpha\log(\alpha)+G_{1}\alpha+\mathcal{O}(\alpha^{3/2-\epsilon})$ (15)
@@
$G_{1}$ $=-\log(e)+\frac{1}{2}\sum_{l=1}^{\infty}2^{-l-1}l\log l+\frac{1}{2}\sum_{a=1}^{\infty}\sum_{b=1}^{\infty}(b+1)2^{-a}2^{-b}h\left(\frac{1}{b+1}\right),$ (16)
which is approximated as $G_{1}\approx 0.4901$.
@@ IV-A Calculation of $H(A^{n},B^{n})$
To calculate $H(A^{n},B^{n}|X^{n},Y,K)$, we use an extended run perspective, where each bit at the two ends of the run is included in the run. Specifically, for a run of 0’s, the last bit of the left neighboring run and the first bit of the right neighboring run are considered part of the extended run; the same applies to runs of 1’s. First, we introduce a modified insertion process, defined by the modified vectors $\hat{A}^{n}$ and $\hat{B}^{n}$. In this modified process, we consider at most a single bit insertion for each extended run. If there is more than one insertion in an extended run, all the insertions in that run (not the extended run) are reversed (i.e., the run is transmitted without any insertion, and the input and output runs are identical). Note that for small insertion probabilities, typical insertion patterns will have insertions that are sufficiently spaced apart and will contribute to the higher-order terms in the capacity. Multiple insertions within a run will not affect the higher-order terms as will be shown later. Hence, one can obtain an accurate estimation of $H(A^{n},B^{n}|X^{n},Y,K)$ when the insertion probability is small (and consequently the capacity).
@@
With the modified insertion process, since we leave the runs unchanged if there are no insertions or exactly one insertion in the corresponding extended run, for these runs, ${\mathbb{Z}}^{n}$ and ${\mathbb{V}}^{n}$ consist entirely of 0’s. For the remaining runs, the corresponding segments in ${\mathbb{Z}}^{n}$ and ${\mathbb{V}}^{n}$ are the same as in ${\mathbb{A}}^{n}$ and ${\mathbb{B}}^{n}$, respectively. Therefore, ${\mathbb{Z}}^{n}$ and ${\mathbb{V}}^{n}$ represent the segments in ${\mathbb{A}}^{n}$ and ${\mathbb{B}}^{n}$ that are reversed due to the modified insertion process, while the segments corresponding to extended runs with at most one insertion remain unchanged.
In the rest of this section, we first derive a bound on the difference between $H(\hat{A}^{n},\hat{B}^{n}|X^{n},\hat{Y},\hat{K})$ and $H({A}^{n},{B}^{n}|X^{n},{Y},{K})$ for the modified insertion process and the original insertion process where $\hat{Y}$ is the output vector when the modified insertion process is applied, and $\hat{K}$ represents the run lengths of the output (defined similarly to (10), including both the bits in the original input and the inserted bits from the modified insertion process). This represents the approximation error incurred by assuming the modified insertion process instead of the original one. Later, we will show that this bound does not affect the first two-order terms, thereby not impacting our capacity approximation. Next, we compute $H(\hat{A}^{n},\hat{B}^{n}|X^{n},\hat{Y},\hat{K})$, which represents the ambiguity in the insertion pattern and inserted bits given the input, output, and associated run vector when the modified insertion process is applied. By combining these two steps, we will arrive at an approximation for $H({A}^{n},{B}^{n}|X^{n},{Y},{K})$.
We begin by examining the difference between $H(\hat{A}^{n},\hat{B}^{n}|X^{n},\hat{Y},\hat{K})$ and $H({A}^{n},{B}^{n}|X^{n},{Y},{K})$ for the modified and original insertion processes considering the following pairs:
@@
$\left((X^{n},Y,K),(X^{n},\hat{Y},\hat{K})\right).$ (25)
Both of these pairs are in the form of $(T,E)$ such that $T$ is a function of $(E,Z^{n},V^{n})$ and $E$ is a function of $(T,Z^{n},V^{n})$. From the chain rule of entropy, we obtain:
@@
Clearly, the processes $\mathbb{Z}$ and $\mathbb{V}$ are stationary, and we have $2H(Z^{n},V^{n})/n\leq 2h(z,v)$. Let us calculate the PMF $P(z=a,v=b)$.
@@
$P(z=1,v=1)$ $=P(z=1)P(v=1|z=1)=\frac{1}{2}P(z=1).$
Without loss of generality, let us focus on the first bit of a run. If there is an insertion at this bit, it will be reversed if at least one more insertion occurs within the extended run containing this run. For simplicity, we can consider the first run (the run containing the bit at position 0), which has a length of $L_{0}$ and extended run length $L_{0}+2$. Conditioned on $L_{0}$, the probability that $\{z=1\}$ is given by $\alpha\left(1-(1-\alpha)^{L_{0}+1}\right)$. Therefore, we can write
$P(z=1)=\alpha\left(1-\mathbb{E}[(1-\alpha)^{L_{0}+1}]\right).$ (35)
@@
$P(z=a,v=b)=\begin{cases}&1-\alpha+\mathbb{E}[\alpha(1-\alpha)^{L_{0}+1}],\\ &\qquad\qquad\text{if }(a,b)=(0,0)\\ &\frac{1}{2}\left(\alpha-\mathbb{E}[\alpha(1-\alpha)^{L_{0}+1}]\right),\\ &\qquad\text{if }(a,b)=(1,0)\text{ or }(1,1)\\ &0,\qquad\qquad\text{otherwise},\end{cases}$ (36)
@@
We further have $(1-\alpha)^{L_{0}+1}=1-\alpha(L_{0}+1)+\mathcal{O}(\alpha^{2})$. Hence, the PMF can be rewritten as:
@@
=$=\begin{cases}1-(\mathbb{E}[L_{0}]+1)\alpha^{2}+\mathcal{O}(\alpha^{3}),&\text{if }(a,b)=(0,0)\\ \frac{1}{2}\mathbb{E}[L_{0}+1]\alpha^{2}+\mathcal{O}(\alpha^{3}),&\text{if }(a,b)=(1,0)\\ &\text{or }(a,b)=(1,1)\\ 0,&\text{otherwise}.\end{cases}$ (37)
One can upper bound the nonzero probabilities in (IV-B) as
$P(z=0,v=0)$ $\leq 1-2\alpha^{2}\mathbb{E}[L_{0}]+\mathcal{O}(\alpha^{3}),$ (38)
$P(z=1,v=0)$ $\leq\alpha^{2}\mathbb{E}[L_{0}]\leq 2\alpha^{2}\mathbb{E}[L_{0}]+\mathcal{O}(\alpha^{3}),$ (39)
$P(z=1,v=1)$ $\leq\alpha^{2}\mathbb{E}[L_{0}]\leq 2\alpha^{2}\mathbb{E}[L_{0}]+\mathcal{O}(\alpha^{3}),$ (40)
which will be used to calculate the dominant terms of $h(z,v)$. This term identifies the loss incurred when restricting ourselves to the modified insertion process, i.e., when we assume there is at most one insertion in an extended run. Intuitively, these probabilities will not contribute to the dominant terms of the capacity expression. This will be made precise by working out the converse.
@@
Consider any $\mathbb{X}$ such that $\mathbb{E}[L_{0}\log L_{0}]<\infty$. Then, we have
@@
where $-2h(z,v)\leq\eta\leq\alpha^{2}\mathbb{E}[L_{0}\log(L_{0})]+2h(z,v)$.
@@
Consider any $\mathbb{X}$ such that $\mathbb{E}[L_{0}\log L_{0}]<\infty$. Then, we have
@@
where $0\leq\delta\leq\alpha^{2}\mathbb{E}[L_{0}\log(L_{0})]$.
@@
The proof consists of the following steps. First, we fix the channel input $x^{n}$ and any possible output $\hat{y}(x^{n})$. Then, we estimate (the $\log$ of) the number of possible realizations of $\hat{A}^{n}$ and $\hat{B}^{n}$ that might lead to the input/output pair $(x^{n},\hat{y})$ given the marker position vector $\hat{k}$. The final step is to take the expectation over $(x^{n},\hat{y},\hat{k})$.
Proceeding from left to right, and using the constraints on both $\hat{\mathbb{A}}^{n}$ and $\hat{\mathbb{B}}^{n}$, one can map each run of $\hat{y}$ to a run of $x^{n}$ given the marker vector $\hat{k}$. To simplify the exposition, let us focus on the $r$-th run of $x^{n}$ which consists of ‘0’s. $l_{r}$ denotes the length of this run (without the extended bits).
@@
    No insertion: Since there is no insertion, there is a one-to-one mapping between the run of $x^{n}$ and $\hat{y}$. Hence, there is no ambiguity and no contribution to $H(\hat{A}^{n},\hat{B}^{n}|x^{n},\hat{y},\hat{k})$.
@@
    Insert bit ‘1’: Since the original run of $x^{n}$ consists of ‘0’s, inserting ‘1’ will reveal all the information regarding $\hat{A}$ and $\hat{B}$ given the marker vector $k$. Hence, there is no contribution to $H(\hat{A}^{n},\hat{B}^{n}|x^{n},\hat{y},\hat{k})$.
@@
$P(\hat{A}_{i}=1)=(1-\alpha)^{l_{(i)}+1}\alpha,$ (47)
and we can write $P(\hat{A}_{i}=1)\in(\alpha-(l_{(i)}+1)\alpha^{2},\alpha)$. Since, $l_{(i)}\geq 1$ and $l_{(i)}+1\leq 2l_{(i)}$, we can also write
$P(\hat{A}_{i}=1)\in(\alpha-2l_{(i)}\alpha^{2},\alpha).$ (48)
@@
where $0\leq\delta\leq\alpha^{2}\mathbb{E}[L_{0}\log(L_{0})]$, which completes the proof of Lemma 6. ∎
@@
For a simple insertion channel model with a stationary and ergodic input process $\mathbb{X}$ satisfying $H(\mathbb{X})>1-\alpha^{\gamma}$, where $\gamma>\frac{1}{2}$, we have:
$H(K|X^{n},Y)=\frac{\alpha}{2}\sum_{a=1}^{\infty}\sum_{b=1}^{\infty}$ $(b+1)2^{-a}2^{-b}h\left(\frac{1}{b+1}\right)+\epsilon_{1},$ (50)
@@
$-\alpha^{1+\gamma/2-\epsilon}-4\alpha^{1+\gamma-\epsilon/2}-\frac{\alpha^{2}}{2}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}(r_{j}+r_{j+1})(r_{j+1}+1)h\left(\frac{1}{r_{j+1}+1}\right)\leq\epsilon_{1}\leq 4\alpha^{1+\gamma-\epsilon/2}+\alpha^{1+\gamma/2-\epsilon},$ (51)
@@
To compute $H(K|X^{n},Y)$, we begin by defining a perturbed insertion process, denoted as $\check{\mathbb{A}}$. For the input process $\mathbb{X}$, we introduce $\check{\mathbb{Z}}$, whose elements correspond to two consecutive input runs with at most one insertion, equated to ‘0’. Otherwise, the elements mirror those of the original insertion process ${\mathbb{A}}$. To clarify further, for two consecutive input runs $\{S_{i},S_{i+1}\}$, we define ˇ$\check{\mathbb{Z}}^{i}$ as a binary process with all zeros when $\{S_{i},S_{i+1}\}$ has at most one insertion. Otherwise, $\check{Z}_{l}^{i}=1$ if $X_{l}\in S_{i}$ and $A_{l}=1$, $\forall i$. Utilizing the definition of $\check{Z}_{l}^{i}$, we further define $\check{\mathbb{Z}}$ in the following way:
@@
$\qquad X(1)X(2)\cdots X(M),\check{Y}(1)\check{Y}(2)\cdots\check{Y}(M)\bigg{)},$ (55)
@@
$|\displaystyle H(\check{K}|X^{n},\check{Y})=\sum_{j=1}^{M}H\bigg{(}|\check{Y}(j)|$$\Big{|}X(j)\cdots X(M),\check{Y}(j)\cdots\check{Y}(M)\bigg{)}.$ (56)
@@
For a simple insertion channel with the perturbed insertion process and a stationary and ergodic input process $\mathbb{X}$ satisfying $H(\mathbb{X})>1-\alpha^{\gamma}$, where $\gamma>\frac{1}{2}$, we have:
$H(\check{K}|X^{n},\check{Y})=\frac{\alpha}{2}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}$ $(r_{j+1}+1)2^{-r_{j}}2^{-r_{j+1}}h\left(\frac{1}{r_{j+1}+1}\right)-\epsilon_{2}+\epsilon_{3},$ (57)
@@
$-\alpha^{1+\gamma/2-\epsilon}\leq\epsilon_{3}\leq\alpha^{1+\gamma/2-\epsilon}.$ (59)
@@
We define the length of $\bar{Y}(j)$ as $L_{j}$, denoted by $|\bar{Y}(j)|=L_{j}$. Consider four cases:
1. When $|\bar{Y}(j)|=|X(j)|$ and $|\bar{Y}(j+1)|=|X(j+1)|$: In this scenario, there is no insertion in $X(j)$. Thus, $t_{j}=0$, since we have sufficient information to determine $|\check{Y}(j)|$, given $X(j)\cdots X(M)$.
2. When $|\bar{Y}(j)|<|X(j)|$: Here, we observe a single-bit insertion inside $X(j)$ with the inserted bit having the opposite polarity, for example, inserting a ‘0’ bit into a run of ‘1’s.
We can directly determine $|\check{Y}(j)|$ using $\bar{Y}(j)$, $\bar{Y}(j+1)$, and $\bar{Y}(j+2)$. Hence, $t_{j}=0$.
Note that $|\bar{Y}(j+1)|=1$, and the corresponding bit has the opposite polarity to $\bar{Y}(j)$ and $\bar{Y}(j+2)$.
3. When $|\bar{Y}(j)|>|X(j)|$ and $|\bar{Y}(j+1)|=|X(j+1)|$: Here, we observe a single-bit insertion inside $X(j)$ with the inserted bit having the same polarity, for example, inserting a ‘0’ bit into a run of ‘0’s.
@@
4. When $|\bar{Y}(j)|=|X(j)|$ and $|\bar{Y}(j+1)|=|X(j+1)|+1$: In this case, a single insertion occurs in $X(j)X(j+1)$, and we cannot distinguish between two different scenarios. Consider the following example:
$X(j)X(j+1)$ =[1 1 1 0 0 0 0]absentdelimited-[]$1$\ 1\ 0\ 0\ 0\ 0]= [ 1 1 1 0 0 0 0 ] (62)
$\bar{Y}(j)\bar{Y}(j+1)$ =[1 1 1 0 0 0 0 0]absentdelimited-[]$1$\ 1\ 0\ 0\ 0\ 0\ 0]= [ 1 1 1 0 0 0 0 0 ] (63)
@@
    $|\check{Y}(j)|$ $=|\bar{Y}(j)|+1=|X(j)|+1$ (64)
    $|\check{Y}(j+1)|$ $=|\bar{Y}(j+1)|-1=|X(j+1)|$ (65)
@@
    $|\check{Y}(j)|$ $=|\bar{Y}(j)|=|X(j)|$ (67)
    $|\check{Y}(j+1)|$ $=|\bar{Y}(j+1)|+1=|X(j+1)|+1$ (68)

### v4

@@ IV Capacity Approximation for the Simple Insertion Channel Model
Let $C_{1}(\alpha)$ be the capacity of the simple insertion channel with insertion probability $\alpha$. Then, for small $\alpha$ and any $\epsilon>0$, we have
$C_{1}(\alpha)=1+\alpha\log(\alpha)+G_{1}\alpha+\mathcal{O}(\alpha^{3/2-\epsilon})$ (14)
@@
$G_{1}$ $=-\log(e)+\frac{1}{2}\sum_{l=1}^{\infty}2^{-l-1}l\log l+\frac{1}{2}\sum_{a=1}^{\infty}(a+1)2^{-a}h\left(\frac{1}{a+1}\right),$ (15)
and it is approximately $0.4901$.
@@ IV-A Calculation of $H(A^{n},B^{n})$
To calculate $H(A^{n},B^{n}|X^{n},Y,K)$, we consider a run perspective approach based on the input run structure. First, we introduce a modified insertion process, defined by the modified vectors $\hat{A}^{n}$ and $\hat{B}^{n}$. In this modified process, we consider at most a single bit insertion for each run. If there is more than one insertion in a run, all the insertions in that run are reversed (i.e., the run is transmitted without any insertion, and the input and output runs are identical). Note that for small insertion probabilities, typical insertion patterns will have insertions that are sufficiently spaced apart and will contribute to the higher-order terms in the capacity. Multiple insertions within a run will not affect the higher-order terms as will be shown later. Hence, one can obtain an accurate estimation of $H(A^{n},B^{n}|X^{n},Y,K)$ when the insertion probability is small (and consequently the capacity).
@@
With the modified insertion process, since we leave the runs unchanged if there are no insertions or exactly one insertion in the corresponding run, for these runs, ${\mathbb{Z}}^{n}$ and ${\mathbb{V}}^{n}$ consist entirely of 0’s. For the remaining runs, the corresponding segments in ${\mathbb{Z}}^{n}$ and ${\mathbb{V}}^{n}$ are the same as in ${\mathbb{A}}^{n}$ and ${\mathbb{B}}^{n}$, respectively. Therefore, ${\mathbb{Z}}^{n}$ and ${\mathbb{V}}^{n}$ represent the segments in ${\mathbb{A}}^{n}$ and ${\mathbb{B}}^{n}$ that are reversed due to the modified insertion process, while the segments corresponding to runs with at most one insertion remain unchanged.
As an example, consider the following binary input vector:
[…]
with insertion pattern and inserted-bit vectors given by
$A^{n}$ =[0 0 0 1 0 1 0 1 $=[0\$0\ $\vdots\$0\ 1\ 0\ 1\ $\vdots\$0\ 1\ 0],
$B^{n}$ =[0 0 0 1 0 0 0 1 $=[0\$0\ $\vdots\$0\ 1\ 0\ 0\ $\vdots\$0\ 1\ 0].
With this insertion pattern, there are two insertions in the second run of $X^{n}$, which means that these insertions are reversed with the modified insertion process, i.e., we have
$\hat{A}^{n}$ =[0 0 0 0 0 0 0 1 $=[0\$0\ $\vdots\$0\ 0\ 0\ 0\ $\vdots\$0\ 1\ 0],
$\hat{B}^{n}$ =[0 0 0 0 0 0 0 1 $=[0\$0\ $\vdots\$0\ 0\ 0\ 0\ $\vdots\$0\ 1\ 0],
and hence the resulting segments of the processes ${\mathbb{Z}}^{n}$ and ${\mathbb{V}}^{n}$ are
$Z^{n}$ =[0 0 0 1 0 1 0 0 $=[0\$0\ $\vdots\$0\ 1\ 0\ 1\ $\vdots\$0\ 0\ 0],
$V^{n}$ =[0 0 0 1 0 0 0 0 $=[0\$0\ $\vdots\$0\ 1\ 0\ 0\ $\vdots\$0\ 0\ 0].
Note that, as already stated in our definition of ${\mathbb{Z}}^{n}$ and ${\mathbb{V}}^{n}$, these segments keep track of the insertions that are reversed by the modified insertion process.
[…]
@@
[…]
Both of these pairs are in the form of $(T,E)$ such that $T$ is a function of $(E,Z^{n},V^{n})$ and $E$ is a function of $(T,Z^{n},V^{n})$. Note that even though there is no exact formulation for $T$ and $E$, for the pair (23), we have $T=(X^{n},Y,K,A^{n},B^{n})$ and $E=(X^{n},\hat{Y},\hat{K},\hat{A}^{n},\hat{B}^{n})$, and $T$ is a function of $(E,Z^{n},V^{n})=(X^{n},\hat{Y},\hat{K},\hat{A}^{n},\hat{B}^{n},Z^{n},V^{n})$; that is, $T$ can be recovered from $(X^{n},\hat{Y},\hat{K},\hat{A}^{n},\hat{B}^{n})$ once we also have the information of $Z^{n}$ and $V^{n}$, which describe the reversed insertions as explained previously. One can extend this analogy in the other direction (i.e., vice versa), as well as to the second pair given in (24).
From the chain rule of entropy, we obtain:
@@
Clearly, the processes $\mathbb{Z}$ and $\mathbb{V}$ are stationary, and we have $2H(Z^{n},V^{n})/n\leq 2h(z,v)$, where $h(z,v)$ denotes the joint entropy of the random variables $z$ and $v$, whose corresponding probabilities are given below. Let us calculate the PMF $P(z=a,v=b)$.
@@
$P(z=1,v=1)$ $=P(z=1)P(v=1|z=1)=\frac{1}{2}P(z=1).$ (34)
Without loss of generality, let us focus on the first bit of a run. If there is an insertion at this bit, it will be reversed if at least one more insertion occurs within the run containing this run. For simplicity, we can consider the first run (the run containing the bit at position 0), which has a length of $L_{0}$. Conditioned on $L_{0}$, the probability that $\{z=1\}$ is given by $\alpha\left(1-(1-\alpha)^{L_{0}-1}\right)$. Therefore, we can write
$P(z=1)=\alpha\left(1-\mathbb{E}[(1-\alpha)^{L_{0}-1}]\right).$ (35)
@@
$P(z=a,v=b)=\begin{cases}&1-\alpha+\mathbb{E}[\alpha(1-\alpha)^{L_{0}-1}],\\ &\hskip 18.49988pt\hskip 18.49988pt\text{if }(a,b)=(0,0)\\ &\frac{1}{2}\left(\alpha-\mathbb{E}[\alpha(1-\alpha)^{L_{0}-1}]\right),\\ &\hskip 18.49988pt\text{if }(a,b)=(1,0)\text{ or }(1,1)\\ &0,\hskip 18.49988pt\hskip 18.49988pt\text{otherwise},\end{cases}$ (36)
@@
We further have $(1-\alpha)^{L_{0}-1}=1-\alpha(L_{0}-1)+\mathcal{O}(\alpha^{2})$. Hence, the PMF can be rewritten as:
@@
$=\begin{cases}1-(\mathbb{E}[L_{0}]-1)\alpha^{2}+\mathcal{O}(\alpha^{3}),&\text{if }(a,b)=(0,0)\\ \frac{1}{2}\mathbb{E}[L_{0}-1]\alpha^{2}+\mathcal{O}(\alpha^{3}),&\text{if }(a,b)=(1,0)\\ &\text{or }(a,b)=(1,1)\\ 0,&\text{otherwise},\end{cases}$ (37)
which will be used to calculate the dominant terms of $h(z,v)$. This term identifies the loss incurred when restricting ourselves to the modified insertion process, i.e., when we assume there is at most one insertion in a run. Intuitively, these probabilities will not contribute to the dominant terms of the capacity expression. This will be made precise by working out the converse.
@@
Consider any $\mathbb{X}\in\mathcal{S}$ such that $\mathbb{E}[L_{0}\log L_{0}]<\infty$. Then, we have
@@
where $2h(z,v)\leq\eta\leq\frac{1}{2}\alpha^{2}\mathbb{E}[L_{0}\log(L_{0})]+2h(z,v)$. Note that $h(z,v)$ can be computed using the probabilities provided in (IV-B). Also, this quantity will be used explicitly in the achievability and converse results later in the paper.
@@
Consider any $\mathbb{X}\in\mathcal{S}$ such that $\mathbb{E}[L_{0}\log L_{0}]<\infty$. Then, we have
@@
where $0\leq\delta\leq\frac{1}{2}\alpha^{2}\mathbb{E}[L_{0}\log(L_{0})]$.
@@
We follow a similar approach to that of [kanoria2010deletion, Lemma IV.4]. The details are given below.
@@
[…]
Proceeding from left to right, and using the constraints on both $\hat{\mathbb{A}}^{n}$ and $\hat{\mathbb{B}}^{n}$, one can map each run of $\hat{y}$ to a run of $x^{n}$ given the marker vector $\hat{k}$. To simplify the exposition, let us focus on the $r$-th run of $x^{n}$ which consists of ‘0’s. $l_{r}$ denotes the length of this run.
@@
[…]
@@
[…]
@@
$P(\hat{A}_{i}=1)$ $=P\left(\text{Having one insertion after the }i\text{-th bit }\right)$
@@
$P(\hat{A}_{i}=1)=(1-\alpha)^{l_{(i)}-1}\alpha,$ (44)
and we can write $P(\hat{A}_{i}=1)\in(\alpha-(l_{(i)}-1)\alpha^{2},\alpha)$, which can be rewritten as
$P(\hat{A}_{i}=1)\in(\alpha-l_{(i)}\alpha^{2},\alpha).$ (45)
Taking the expectation of (42), we obtain
$\frac{1}{2}\sum_{i=1}^{n}\left(1\cdot P(\hat{A}_{i}=1)+0\cdot P(\hat{A}_{i}=0)\right)\mathbb{E}[\log(l_{(i)})]=\frac{1}{2}\sum_{i=1}^{n}P(\hat{A}_{i}=1)\mathbb{E}[\log(l_{(i)})].$ (46)
Letting $n\rightarrow\infty$ and using the bound given in (45), we can equivalently write
@@
where $0\leq\delta\leq\frac{1}{2}\alpha^{2}\mathbb{E}[L_{0}\log(L_{0})]$, which completes the proof of Lemma 6.
@@
For a simple insertion channel with any $\epsilon>0$, there exists $\alpha_{0}\equiv\alpha_{0}(\epsilon)>0$ such that for any $\alpha<\alpha_{0}$ and any stationary and ergodic input process $\mathbb{X}\in\mathcal{S}_{\lfloor 1/\alpha\rfloor}$ satisfying $H(\mathbb{X})>1-\alpha^{\gamma}$, where $\gamma>\frac{1}{2}$, we have:
$\lim_{n\rightarrow\infty}\frac{1}{n}H(K|X^{n},Y)=\frac{\alpha}{2}\sum_{a=1}^{\infty}(a+1)2^{-a}h\left(\frac{1}{a+1}\right)+\epsilon_{1},$ (48)
@@
$\alpha^{1+\gamma/2-\epsilon}-4\alpha^{1+\gamma-\epsilon/2}-\frac{\alpha^{2}}{2}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}(r_{j}+r_{j+1})(r_{j+1}+1)h\left(\frac{1}{r_{j+1}+1}\right)\leq\epsilon_{1}\leq 4\alpha^{1+\gamma-\epsilon/2}+\alpha^{1+\gamma/2-\epsilon},$ (49)
@@
[…]
@@
$\hskip 18.49988ptX(1)X(2)\cdots X(M),\check{Y}(1)\check{Y}(2)\cdots\check{Y}(M)\bigg),$ (53)
@@
$H(\check{K}|X^{n},\check{Y})=\sum_{j=1}^{M}H\bigg(|\check{Y}(j)|$ $\Big|X(j)\cdots X(M),\check{Y}(j)\cdots\check{Y}(M)\bigg).$ (54)
@@
For a simple insertion channel using the perturbed insertion process with any $\epsilon>0$, there exists $\alpha_{0}\equiv\alpha_{0}(\epsilon)>0$ such that for any $\alpha<\alpha_{0}$ and any stationary and ergodic input process $\mathbb{X}\in\mathcal{S}_{\lfloor 1/\alpha\rfloor}$ satisfying $H(\mathbb{X})>1-\alpha^{\gamma}$, where $\gamma>\frac{1}{2}$, we have:
$\lim_{n\rightarrow\infty}\frac{1}{n}H(\check{K}|X^{n},\check{Y})=\frac{\alpha}{2}\sum_{r_{j+1}=1}^{\infty}(r_{j+1}+1)2^{-r_{j+1}}h\left(\frac{1}{r_{j+1}+1}\right)-\epsilon_{2}+\epsilon_{3},$ (55)
@@
$\alpha^{1+\gamma/2-\epsilon}\leq\epsilon_{3}\leq\alpha^{1+\gamma/2-\epsilon}.$ (57)
@@
We note that our procedure proceeds from left to right. Since insertions may introduce new runs under our definition of $\bar{Y}$, the indices of $\check{Y}(j)$ and $\bar{Y}(j^{\prime})$ may not coincide, even though the corresponding runs themselves do match. However, by starting from the left, one can cancel the runs that are resolved on that side and then continue the process accordingly. For example, the first bit of $\check{Y}(j)$ will correspond to the first bit of $\bar{Y}(j^{\prime})$; hence, although there may be an index mismatch between $j$ and $j^{\prime}$, the bits themselves are aligned and represent the same run segment.
Consider four cases:
1. When $|\bar{Y}(j^{\prime})|=|X(j)|$ and $|\bar{Y}(j^{\prime}+1)|=|X(j+1)|$: In this scenario, there is no insertion in $X(j)$. Thus, $t_{j}=0$, since we have sufficient information to determine $|\check{Y}(j)|$, given $X(j)\cdots X(M)$.
2. When $|\bar{Y}(j^{\prime})|<|X(j)|$: Here, we observe a single-bit insertion inside $X(j)$ with the inserted bit having the opposite polarity, for example, inserting a ‘0’ bit into a run of ‘1’s.
We can directly determine $|\check{Y}(j)|$ using $\bar{Y}(j^{\prime})$, $\bar{Y}(j^{\prime}+1)$, and $\bar{Y}(j^{\prime}+2)$. Hence, $t_{j}=0$.
Note that $|\bar{Y}(j^{\prime}+1)|=1$, and the corresponding bit has the opposite polarity to $\bar{Y}(j^{\prime})$ and $\bar{Y}(j^{\prime}+2)$.
3. When $|\bar{Y}(j^{\prime})|>|X(j)|$ and $|\bar{Y}(j^{\prime}+1)|=|X(j+1)|$: Here, we observe a single-bit insertion inside $X(j)$ with the inserted bit having the same polarity, for example, inserting a ‘0’ bit into a run of ‘0’s.
@@
4. When $|\bar{Y}(j^{\prime})|=|X(j)|$ and $|\bar{Y}(j^{\prime}+1)|=|X(j+1)|+1$: In this case, a single insertion occurs in $X(j)X(j+1)$, and we cannot distinguish between two different scenarios. Consider the following example:
$X(j)X(j+1)$ =[1 1 $0 0]\displaystyle=[1\$1\ 1\ 0\ 0\ 0\ 0] (60)
$\bar{Y}(j^{\prime})\bar{Y}(j^{\prime}+1)$ $=[1\ 1\ 1\ 0\ 0\ 0\ 0\ 0]$ (61)
@@
    $|\check{Y}(j)|$ $=|\bar{Y}(j^{\prime})|+1=|X(j)|+1$ (62)
    $|\check{Y}(j+1)|$ $=|\bar{Y}(j^{\prime}+1)|-1=|X(j+1)|$ (63)
@@
    $|\check{Y}(j)|$ $=|\bar{Y}(j^{\prime})|=|X(j)|$ (65)
    $|\check{Y}(j+1)|$ $=|\bar{Y}(j^{\prime}+1)|+1=|X(j+1)|+1$ (66)